# Perturbation-based Effect Measures for Compositional Data

### v2

@@ 1 Introduction
Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982), which we discuss in more detail in Section 1.1. A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates. More recently, in the context of high-dimensional and sparse data, alternative approaches (Knight et al., 2018; Cammarota et al., 2020; Huang et al., 2023) to solving the second goal have been proposed that apply flexible machine learning methods to model the relationship between $Y$ and $Z$ nonparametrically. However, such approaches are at odds with the first goal of providing interpretable effect measures and cannot be combined with the above-mentioned approaches which heavily rely on parametric models to define such effects.
@@
Our main contribution is to provide a unified framework for defining interpretable effect measures for compositional features which we call average perturbation effects that can be combined with flexible nonparametric estimation procedures. Special cases of these average perturbation effects have interpretations that are similar to some model-based individual feature effect measures, e.g., Cox-directions (Cox, 1971) and log-contrast coefficients (Aitchison and Bacon-Shone, 1984). Average perturbation effects, however, are defined without reference to a specific regression model and therefore do not rely on a correctly specified parametric model. Importantly, this allows us to construct more flexible target quantities, two of which we believe are particularly relevant for applications. First, effects based on binary perturbations are useful when summarizing the effect of a discrete change, e.g. the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}$ to the simplex. In Section 4.3 we show empirically that the compositional features with significant effects as measured by these binary perturbations coincide with those deemed important by the standard log-contrast regression procedures used in the microbiome literature. Second, effects based on directional perturbations summarize continuous changes in specific directions. This can involve changing individual components of $Z$, but also allows for more complex changes that are not tied to a specific coordinate of $Z$. For example, we can define the effect of ‘diversifying’ $Z$, i.e. moving closer to the center of $\Delta^{d-1}$. Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity. However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Example 2 and Section 4.2). In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
@@
Average perturbation effects are defined by comparing the change in the expected response of $Y$ under a hypothetical perturbation of the composition $Z$. To motivate this, consider Euclidean covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$. Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function. Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the rate of change in the $j$th coordinate, or more formally,
@@
Therefore, the coefficient $\beta^{j}$ also corresponds to the initial rate of change (initial referring to $\gamma=0$) of $f$ along the perturbation $\psi$. Similar observations are possible regarding other interpretable effect measures for Euclidean covariates and under more general nonlinear models.
@@
where we implicitly assume $\mathbb{P}(T=0)>0$ and $\mathbb{P}(T=1)>0$. It is well-known that this type of marginal dependence analysis can be misleading if dependencies exist between the individual components. [^1] This is true when $Z$ is Euclidean, but the problem is even more severe when $Z$ is compositional as the sum-to-one constraint enforces additional dependency between the components of $Z$. To illustrate this, assume the disease indicator given the composition has conditional distribution for all $z\in\Delta^{d-1}$ given by
$Y\,|\,Z=z\sim\operatorname{Ber}(\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{z^{2}=0\}}),$
that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ decreases the probability of having the disease. In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1. If $Z$ was Euclidean, we could adjust for the dependence in the remaining coordinates by estimating a conditional treatment effect such as
@@
and marginalizing it to produce a single effect measure $\mathbb{E}[h(Z^{-j})]$. [^2] In the literature on causal inference (and under appropriate causal assumptions), $h$ corresponds to the conditional average treatment effect (CATE). Estimating this quantity is a well-studied problem, e.g. using the AIPW estimator (Robins et al., 1994) or using a partially linear model (Chernozhukov et al., 2018). Unfortunately, neither of these estimators are well-defined when $Z$ is compositional as it is not possible to change $T$ and $Z^{-j}$ independently of each other and, in fact, for all $z^{-j}$, we have $h(z^{-j})=0$ as $Y\mbox{${}\perp\mkern-11.0mu\perp{}$}T\,|\,Z^{-j}$ due to the sum-to-one constraint which makes $T$ fully determined by $Z^{-j}$. In the Table 1, we provide the results from applying our proposed marginalized conditional treatment effect the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
@@
Table 1: Numerical illustration of Examples 1 and 2. The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model. The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators. Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect. (Left) The estimated MTE is negative, even though the true effect is positive. The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign. (Right) The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction. The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect. Details on both experiments are provided in Section S4.1 of the supplementary material.
@@
Consider the problem of determining the effect of an increase in diversity of $Z$ on a response $Y$. Such a question is of immediate interest in many applied fields, e.g. when investigating whether more racially diverse university programs improve educational outcomes (Antonio et al., 2004) or whether increased income equality boosts economic output (Shin, 2012). The prevalent approach is to summarize diversity via a one-dimensional summary statistic $G(Z)$ (e.g., the Gini coefficient, generalized entropy or Gini-Simpson index) and analyze the correlation between $G(Z)$ and $Y$. As in Example 1 this type of marginal dependence analysis is potentially misleading in practice. In Table 1 we illustrate this on a specific numeric example, where the true causal effect is positive, but the marginal method leads to a negative effect of diversity (details are provided in Section S4.1 of the supplementary material). To avoid this type of confounding bias, one would again like to adjust for the remaining dependence within $Z$. However, since $G(Z)$ depends on all of $Z$ existing adjustment approaches are not applicable. The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjusts for the confounding as shown in Table 1. A more in-depth example in which the marginal approach is misleading is given in Section 4.2.1 where we estimate the effect of racial diversity on income on a semi-synthetic dataset based on US census data. A real application is given in Section 4.2.2 where the effect of racial diversity on school grades is analyzed using data from schools in New York. Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained Euclidean data. Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
@@
We postpone the formal definitions of perturbations and average perturbation effects to Section 2 but give a brief description here. A perturbation is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed. For example, in Example 1 when considering the first microbe (i.e., $j=1$), we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one. By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$. The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual Euclidean approach outlined in Example 1. By marginalizing over $Z$ the perturbation-based approach leads to the average effect
@@ 1.1 Related literature
Model-based effect estimation with compositional variables has a long history and was perhaps first thoroughly investigated under the name ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002). While parts of the literature focused on the problem of selecting design points $z_{1},\dots,z_{n}\in\Delta^{d-1}$ in optimal ways, a key aspect was also the problem of constructing reasonable regression models in this setting. Scheffé (1958) pointed out that polynomial models are overidentified with compositional data due to the sum-to-one constraint and provided a canonical choice of polynomials. Becker (1968) critiqued the use of polynomials for compositional features and instead provided parametric models with homogenous functions of degree one as building blocks. Cox (1971) argued that the polynomials of Scheffé would often not result in interpretable coefficients but that one should instead define the coefficients differently, e.g. so that the $j$th coefficient in a linear model corresponds to increasing the value of $Z^{j}$ while scaling the remaining components towards a fixed and problem-dependent ‘standard mixture’ $z_{*}\in\Delta^{d-1}$. It is possible to define a model-agnostic version of Cox’s linear effects using our perturbation-based framework (see Proposition S4 in Section S6 of the supplementary material). Draper and John (1977) developed predictive polynomial models based on the reciprocal of the components of $Z$ to cope with cases where the expected response diverges at the boundary. Aitchison and Bacon-Shone (1984) introduced the log-contrast regression model where the linear version corresponds to a linear regression on the log-transformed $Z$ with a sum-to-zero constraint on the coefficients. Log-contrast models have seen extensive use in practice and have since been generalized to high-dimensional $Z$ (Lin et al., 2014; Simpson et al., 2021). One of our proposed feature influence measures, the compositional feature influence with multiplicative speed (see Example 3), corresponds exactly to the log-contrast coefficients when the model is correctly specified (see Proposition S3 in Section S6 of the supplementary material or Huang et al. (2023) where this was first shown). Defining practical predictive models for compositional data remains an active topic of research (Xiong et al., 2015; Fiksel et al., 2022).
@@
Naturally, our proposed average perturbation effects are also connected to the broader literature on compositional data analysis. It is well-known that many conventional statistical procedures become invalid when applied to compositional data. For example, already Pearson (1897) drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint. The most prevalent class of general methods is based on the seminal paper by Aitchison (1982) which introduces the log-ratio approach. In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms (Aitchison, 1982, 1986; Egozcue et al., 2003) and then performing the statistical analysis in the transformed space. Our approach is not based on log-ratios, but it is possible to construct perturbations from this perspective (see Section 2.1.3). The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however this in principle requires $Z$ to take values in the open simplex. This is known as the zero problem and has sparked a large body of research developing various types of zero-imputation schemes (e.g., Martín-Fernández et al., 2000, 2003, 2015; Palarea-Albaladejo and Martín-Fernández, 2008; Lubbe et al., 2021) that can be used before applying the log-ratio approach. Greenacre et al. (2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches. While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa.
@@
There are other approaches to the analysis of data on the simplex, both in the form of alternative transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022), or by appealing to more general notions of object-oriented data analysis (Marron and Dryden, 2021). There has also been much recent work on manifold-valued data where compositional data is a special case (Dai and Müller, 2018; Dai et al., 2021; Petersen et al., 2022).

### v3

@@ 1 Introduction
Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z=(Z^{1},\dots,Z^{d})\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982). We briefly summarize the developments of these fields in Section 1.1. A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates. More recently, in the context of high-dimensional and sparse data, alternative approaches (Knight et al., 2018; Cammarota et al., 2020; Huang et al., 2023) to solving the second goal have been proposed that apply flexible machine learning methods to model the relationship between $Y$ and $Z$ nonparametrically. However, such approaches are at odds with the first goal of providing interpretable effect measures and cannot be combined with the above-mentioned approaches which heavily rely on parametric models to define such effects.
@@
Our main contribution is to provide a unified framework for defining interpretable effect measures for compositional features which we call average perturbation effects that can be combined with flexible nonparametric estimation procedures. Special cases of these average perturbation effects have interpretations that are similar to some model-based individual feature effect measures, e.g., Cox-directions (Cox, 1971) and log-contrast coefficients (Aitchison and Bacon-Shone, 1984). Average perturbation effects, however, are defined without reference to a specific regression model and therefore do not rely on a correctly specified parametric model. Importantly, this allows us to construct more flexible target quantities, two of which we believe are particularly relevant for applications. First, effects based on binary perturbations are useful when summarizing the effect of a discrete change, e.g. the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}:=(Z^{1},\dots,Z^{j-1},Z^{j+1},\dots,Z^{d})$ to the simplex. In Section 4.2 we show empirically that the compositional features with significant effects as measured by these binary perturbations coincide with those deemed important by the standard log-contrast regression procedures used in the microbiome literature. Second, effects based on directional perturbations summarize continuous changes in specific directions. This can involve changing individual components of $Z$, but also allows for more complex changes that are not tied to a specific coordinate of $Z$. For example, we can define the effect of ‘diversifying’ $Z$, i.e. moving closer to the center of $\Delta^{d-1}$. Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity. However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Example 2 and Section 4.1). In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
@@
Average perturbation effects are defined by comparing the change in the expected response of $Y$ under a hypothetical perturbation of the composition $Z$. To motivate this, consider unconstrained covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$. Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function. Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally,
@@
Therefore, the coefficient $\beta^{j}$ also corresponds to the initial rate of change (initial referring to $\gamma=0$) of $f$ along the perturbation $\psi$. Similar observations are possible regarding other interpretable effect measures for unconstrained covariates and under more general nonlinear models.
@@
where we implicitly assume $\mathbb{P}(T=0)>0$ and $\mathbb{P}(T=1)>0$. It is well-known that this type of marginal dependence analysis can be misleading if dependencies exist between the individual components. [^1] This is true when $Z\in\mathbb{R}^{d}$ is unconstrained, but the problem is even more severe when $Z$ is compositional as the sum-to-one constraint enforces additional dependency between the components of $Z$. To illustrate this, assume the disease indicator $Y$ given the composition $Z$ has conditional distribution for all $z\in\Delta^{d-1}$ given by a Bernoulli distribution with
$\mathbb{P}(Y=1\,|\,Z=z)=\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{z^{2}=0\}}$
that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ reduces the probability of having the disease. In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1 (details are provided in Section S4.1 of the supplementary material). If $Z\in\mathbb{R}^{d}$ was unconstrained, we could adjust for the dependence in the remaining coordinates by estimating a conditional treatment effect such as
@@
and marginalizing it to produce a single effect measure $\mathbb{E}[h(Z^{-j})]$. [^2] In the literature on causal inference (and under appropriate causal assumptions), $h$ corresponds to the conditional average treatment effect (CATE). Estimating this quantity is a well-studied problem, e.g. using the AIPW estimator (Robins et al., 1994) or using a partially linear model (Chernozhukov et al., 2018). Unfortunately, neither of the estimators is well defined when $Z\in\Delta^{d-1}$ as it is not possible to change $T$ and $Z^{-j}$ independently of each other and, in fact, for all $z^{-j}$, we have $h(z^{-j})=0$ as $Y$ is conditionally independent of $T$ given $Z^{-j}$ due to the sum-to-one constraint which makes $T$ fully determined by $Z^{-j}$. In the Table 1, we provide the results from applying our proposed marginalized conditional treatment effect the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
@@
Table 1: Numerical illustration of Examples 1 and 2. The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model. The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators. Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect. Left: The estimated MTE is negative, even though the true effect is positive. The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign. Right: The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction. The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect. Details on both experiments are provided in Section S4.1 of the supplementary material.
@@
Consider the problem of determining the effect of an increase in diversity of $Z$ on a response $Y$. Such a question is of immediate interest in many applied fields, e.g. when investigating whether more racially diverse university programs improve educational outcomes (Antonio et al., 2004) or whether increased income equality boosts economic output (Shin, 2012). The prevalent approach is to summarize diversity via a one-dimensional summary statistic $G(Z)$ (e.g., the Gini coefficient, generalized entropy or Gini-Simpson index) and analyze the correlation between $G(Z)$ and $Y$. As in Example 1 this type of marginal dependence analysis is potentially misleading in practice. In Table 1 we illustrate this on a specific numeric example, where the true causal effect is positive, but the marginal method leads to a negative effect of diversity (details are provided in Section S4.1 of the supplementary material). To avoid this type of confounding bias, one would again like to adjust for the remaining dependence within $Z$. However, since $G(Z)$ depends on all of $Z$ existing adjustment approaches are not applicable. The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjusts for the confounding as shown in Table 1. A more in-depth example in which the marginal approach is misleading is given in Section 4.1.1 where we estimate the effect of racial diversity on income on a semi-synthetic dataset based on US census data. A real application is given in Section 4.1.2 where the effect of racial diversity on school grades is analyzed using data from schools in New York. Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained data. Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
@@
We postpone the formal definitions of perturbations and average perturbation effects to Section 2 but give a brief description here. A perturbation is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed. For example, in Example 1 when considering the first microbe (i.e., $j=1$), we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one. By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$. The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual unconstrained approach outlined in Example 1. By marginalizing over $Z$ the perturbation-based approach leads to the average effect
@@ 1.1 Related literature
Model-based effect estimation with compositional variables has a long history and was perhaps first thoroughly investigated under the name ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002). While parts of the literature focused on the problem of selecting design points $z_{1},\dots,z_{n}\in\Delta^{d-1}$ in optimal ways, a key aspect was also the problem of constructing reasonable regression models in this setting. Scheffé (1958) pointed out that polynomial models are overidentified with compositional data due to the sum-to-one constraint and provided a canonical choice of polynomials. Becker (1968) critiqued the use of polynomials for compositional features and instead provided parametric models with homogenous functions of degree one as building blocks. Cox (1971) argued that the polynomials of Scheffé would often not result in interpretable coefficients but that one should instead define the coefficients differently, e.g. so that the $j$th coefficient in a linear model corresponds to increasing the value of $Z^{j}$ while scaling the remaining components towards a fixed and problem-dependent ‘standard mixture’ $z_{*}\in\Delta^{d-1}$. It is possible to define a model-agnostic version of Cox’s linear effects using our perturbation-based framework (see Proposition S4 in Section S6 of the supplementary material). Draper and John (1977) developed predictive polynomial models based on the reciprocal of the components of $Z$ to cope with cases where the expected response diverges at the boundary. Aitchison and Bacon-Shone (1984) introduced the log-contrast regression model where the linear version corresponds to a linear regression on the log-transformed $Z$ with a sum-to-zero constraint on the coefficients. Log-contrast models have seen extensive use in practice and have since been generalized to high-dimensional $Z$ (Lin et al., 2014; Simpson et al., 2021). One of our proposed feature influence measures, the compositional feature influence with multiplicative speed (see Example 3), corresponds exactly to the log-contrast coefficients when the model is correctly specified (see Proposition S3 in Section S6 of the supplementary material or Huang et al. (2023) where this was first shown). Defining practical predictive models for compositional data remains an active topic of research (Xiong et al., 2015; Fiksel et al., 2022; Li et al., 2023; Bhaduri et al., 2025).
@@
Naturally, our proposed average perturbation effects are also connected to the broader literature on compositional data analysis. It is well-known that many conventional statistical procedures become invalid when applied to compositional data. For example, already Pearson (1897) drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint. The most prevalent class of general methods is based on the seminal paper by Aitchison (1982) which introduces the log-ratio approach. In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms (Aitchison, 1982, 1986; Egozcue et al., 2003) and then performing the statistical analysis in the transformed space. Our approach is not based on log-ratios, but it is possible to construct perturbations from this perspective (see Section 2.1.3). The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however this in principle requires $Z$ to take values in the open simplex. This is known as the zero problem and has sparked a large body of research developing various types of zero-imputation schemes (e.g., Martín-Fernández et al., 2000, 2003, 2015; Palarea-Albaladejo and Martín-Fernández, 2008; Lubbe et al., 2021) that can be used before applying the log-ratio approach. Greenacre et al. (2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches which are further highlighted in Scealy and Welsh (2014). While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa.
@@
There are other approaches to the analysis of data on the simplex, both in the form of alternative transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022), or by appealing to more general notions of object-oriented data analysis (Marron and Dryden, 2021). There has also been much recent work on manifold-valued data where compositional data is a special case (Dai and Müller, 2018; Dai et al., 2021; Petersen et al., 2022; Kurisu et al., 2024).